Formula If you know that n is 1/3 of m.
To find : How could you rewrite this formula?
P=9n-3m

Answer by Theo(13342)

equation is p = 9 * n - 3 * m

you are given that n = 1/3 * m.

replace n with 1/3 * m to get:

p = 9 * 1/3 * m - 3 * m

simplify to get:

p = 3 * m - 3 * m

solve for p to get:

p = 0